# Symmetry verification for noisy quantum simulations of non-Abelian lattice gauge theories

### v1

@@ Abstract
Non-Abelian gauge theories underlie our understanding of fundamental forces of modern physics. Simulating them on quantum hardware is an outstanding challenge in the rapidly evolving field of quantum simulation. A key prerequisite is the protection of local gauge symmetries against errors that, if unchecked, would lead to unphysical results. While an extensive toolkit devoted to identifying, mitigating, and ultimately correcting such errors has been developed for Abelian groups, non-commuting symmetry operators complicate the implementation of similar schemes in non-Abelian theories. Here, we discuss two techniques for error mitigation through symmetry verification, tailored for non-Abelian lattice gauge theories implemented in noisy qudit hardware: dynamical post-selection (DPS), based on mid-circuit measurements without active feedback, and post-processed symmetry verification (PSV), which combines measurements of correlations between target observables and gauge transformations. We illustrate both approaches for the discrete non-Abelian group $D_{3}$ in 2+1 dimensions, illustrating their usefulness for current NISQ devices even in presence of fast fluctuating noise. Our results open new avenues for robust quantum simulation of non-Abelian gauge theories, for further development of error-mitigation techniques, and for measurement-based control methods in qudit platforms.
I Introduction
@@
The rest of the paper is organized as follows: To make this work self-contained, in Sec. II we introduce the formalism needed to implement a quantum simulation of a non-Abelian LGT on qudit devices. In Sec. III, we explain the general idea of the two techniques for ensuring non-Abelian gauge invariance. Section IV specializes the concepts introduced so far to the $D_{3}$ model and presents our numerical analysis of the two protection schemes. Finally, we provide an overview of our work and discuss possible developments in Sec. V. Further numerical results and details are presented in the Appendix.
II Hamiltonian real-time evolution for discrete non-Abelian LGTs
@@
In this work, we are interested in a digital (Totterized) time evolution of LGT models with a discrete non-Abelian symmetry group. While much of our concepts are general, we will focus on a pure gauge theory in 2+1 space-time dimensions in the Kogut–Susskind Hamiltonian formulation. We consider a square lattice with gauge variables residing on its links, see Fig. 1. To provide the necessary background, we review the basic formalism for quantum simulations of LGT models with a discrete non-Abelian group $G$, following the notation of Refs. [53, 54].
@@
Equation (II) can be verified by combining Eq. (II) with the representation of $U^{j}_{mn}$ in the computational basis {|$\{\ket{g}\}$.
@@
where $n_{t}=t/\Delta t$ is the number of Trotter steps. To efficiently encode groups of order $|G|>2$, we specialize our discussion to universal qudit devices  [36, 37, 38, 39, 40], where each fundamental unit of quantum information possesses more than two levels ($d>2$). This enables the implementation of physically local terms using single qudit gates  [54, 58, 59]. The general framework developed for LGTs on qubit digital quantum simulators [60, 61, 54, 62] has already been expanded for qudit quantum hardware [63, 59, 18, 64, 65, 66], showing promising potential for future advances in this direction. See appendix A for a summary of the essential ingredients. For an efficient implementation, we assume the number of levels in the qudit to be at least as large as the group order, $d\geq|G|$.
III Verification of non-Abelian gauge symmetries
@@
III.1 Dynamical post-selection
@@
To further understand the result of this procedure, let us consider a situation where the incoherent error channel in Eq. (9) is described by probabilistically occurring unitaries. Such a scenario includes a broad range of errors, such as uncontrolled gate fluctuations, and will be at the basis of the noise model we use in our numerics in Sec. IV. Assuming the state at time $t$ satisfied the non-Abelian Gauss’ laws, we can approximate a noisy Trotter step of a single random (unitary) trajectory as
@@
III.2 Post-processed symmetry verification
@@
In the case of a lattice gauge theory with a finite (Abelian or non-Abelian) symmetry group $G$, the symmetry-preserving subspace S is given by invariance under the gauge transformations $\Theta_{g,v}\ket{\psi}=\ket{\psi}$ for all group elements $g\in G$ and all of the $n_{\mathrm{v}}$ vertices $v\in\mathrm{V}$. The projector $\Pi_{\mathrm{s}}$ thus takes the form
@@
While this approach does not suffer from an exponentially decreasing number of valid shots, as does usual direct post-selection, the decompositions of Eq. (19) employ on the order of $|G|^{n_{\mathrm{v}}}$ operators. In Sec. IV.2.2, we discuss the resource requirements in more detail. Importantly, implementing this procedure does not require additional circuit elements in the form of ancillas or measurement gates and is not dependent on the commutation relations of the symmetry transformations.
IV Numerical benchmarks on the $D_{3}$ model
@@
We numerically simulate the system dynamics with exact diagonalization. In our simulations, we use a fixed Trotter step $\Delta t/a=0.25$ and different noise strengths $\gamma$, and we set $\frac{1}{g^{2}}=0.5$ in order to work in a scenario where $H_{E}$ and $H_{B}$ compete.
IV.1 Dynamical post-selection
@@
Figure 4: (a) Data clouds represent $\langle\mathcal{O}_{P1}\rangle$ at time $t/a$ on individual trajectories for DPS (red) and noisy (green) evolutions. The opacity reflects the fraction of trajectories retained after DPS. Solid lines represent ensemble averages. The DPS ensemble average follows very well the error-free evolution (black dots). In the inset, we show the standard deviation of DPS data (red dots) and the fraction $P_{\textrm{sur}}$ of trajectories that always preserved the gauge symmetries (black dots). (b-c) Gauge violations on vertex $v_{1}$ for a single successful DPS trajectory (b) and for the average of noisy evolutions without DPS (c). Reflection terms $g=3,4,5$ (solid lines) do not superimpose on a single noisy trajectory, while their averages do. Rotations are here represented by $g=2$ (dashed lines). Notably, the violation of a symmetry $\Theta_{g,v}$ can be partially suppressed also when other gauge charges are measured. The system is evolved with $\gamma=0.2$, $0.51/g^{2}=$ 0.5, and $\Delta t/a=0.25$ for $M=5000$ trajectories.
@@
Here, $k_{g}$ is a normalization factor depending on $g$ such that $\textrm{GV}\in[0,1]$. The Gauss’ laws violation allows us to qualitatively compare post-selected and noisy dynamics and to assess the quality of the gauge-protection protocol.
Figure 4(b) displays the dynamics of $\textrm{GV}_{g,v1}$ at the first vertex for a trajectory where the local symmetries are always successfully protected. Out of the six possible gauge transformations for each vertex, we can neglect the identity $g=0$ as it acts trivially. Moreover, we only monitor the violations corresponding to one rotation ($g=2$), since the $\mathbb{Z}_{3}$ subgroup of $\frac{2\pi}{3}$ rotations is cyclic, rendering conservation of the associated local symmetries equivalent. As shown in Fig. 4(b), the DPS protocol repeatedly suppresses the gauge violations through the auxiliary measurements, and their growth between two successive measurements of the same symmetry operator remains limited. Moreover, gauge violations in Fig. 4(b) decrease even when other symmetries are monitored, which indicates that gauge violations of different local charges are correlated. The behaviour is markedly different in the bare noisy dynamics, shown in Fig. 4(c). Here, the ensemble-averaged violations increase until they reach a plateau, $0.5$ for reflections and $1/\sqrt{3}$ for rotations, corresponding to a fully mixed state [^2] The precise asymptotic value depends on the definition in Eq. (23), but in general every charge sector is uniformly populated after a characteristic time depending on $\gamma$..
@@
IV.2 Comparing DPS and PSV
Let us now explicitly compare the error-mitigation capacity of the two protocols described in Sec. III. First, we contrast the protection efficacy of both methods. After that, we discuss the relative scalings, which helps in assessing which method may be most suitable for a given platform.
IV.2.1 Protection efficacy
@@ IV.2.2 Resource requirements
Additionally, one can estimate Eq. (19) through a random sampling of the group element combinations $\{g_{1},\dots,g_{n_{\mathrm{v}}}\}$. For every sample, one simultaneously measures the prescribed gauge transformation operator and its correlator with the gauge-invariant target observable (or cardinality-weighted cliques thereof). Following analogous considerations for the qubit case [48, 47, 68], one may expect this approach to require a sampling overhead scaling polynomially in $\mathrm{Tr}[\Pi_{\mathrm{s}}\rho]$. The decrease of $\mathrm{Tr}[\Pi_{\mathrm{s}}\rho]$ with circuit depth and system size is governed by the strength of (gauge-invariance breaking) error rates, and therefore the required sampling cost immediately benefits from any quantum hardware developments. In particular, for a given shot budget and error rate, one has a window of simulated time within which controlled sampling of the gauge-invariant expectation value is possible.
V Conclusions
@@
DPS is remarkably resilient against noise and the measurement strategy can be easily adapted to suit specific needs. For instance, the density of mid-circuit measurements can be tuned according to the system studied and the error channel expected on a specific hardware. The only requirement is that the measurement frequency must be above the quantum Zeno threshold for error mitigation in the system, otherwise, the system would enter an unprotected regime with random dynamics jumping between different gauge sectors [41]. The two-qudit gates used in the DPS circuit also contribute with an additional error rate. Nevertheless, as long as the number of additional two-qudit gates is sufficiently smaller than the ones already required for the Hamiltonian time evolution, this does not constitute a decisive limiting factor. For example, in the model we studied and on a qudit-based architecture, one needs $60$ controlled permutations per Trotter step, whereas checking one local reflection (rotation) requires $8$ ($10$) controlled permutations. This advantage shows a promising window of opportunity for future experimental studies on qudit-based hardware during the NISQ era. Moreover, the resource requirements can be mitigated by choosing an optimal subset of group elements for the symmetry measurements: gauge violations on a single vertex $v$ can be fully suppressed by monitoring only a suitably chosen generating set of $G$.
@@ Appendix A Trotterized time evolution of a discrete LGT
The unitary operator generated by the four-body interaction described by the magnetic term $H_{B}$ in Eq. (5) is more complicated to implement. A possible circuit decomposition [54], depicted in Fig. 6, is a generalization of a standard four-body Pauli-$Z$ interaction for qubits: three qudit controlled-multiplication gates acting on the fourth, onto which subsequently a trace gate is applied. Finally, the state of the four qudit registers is restored by applying the inverse of the multiplication gates. The procedure also requires the application of the inversion gate, which maps an element of the group $g$ into its inverse $\mathfrak{U}_{-1}\ket{g}=\ket{g^{-1}}$. The trace gate is diagonal in the group-element basis and implements the parametric operation
@@
The controlled-multiplication gate is defined as
@@
where the first qudit is the control and the second is the target. This is the most costly operation in quantum hardware, as it applies $|G|$ different operations (one of which, the identity, is trivial) depending on the state of the control qudit. The controlled-multiplication (or controlled-permutation) gate $\mathfrak{U}_{X}$ is the only two-qudit unitary required for the time evolution.

### v2

@@ Abstract
Non-Abelian gauge theories underlie our understanding of fundamental forces of modern physics. Simulating them on quantum hardware is an outstanding challenge in the rapidly evolving field of quantum simulation. A key prerequisite is the protection of local gauge symmetries against errors that, if unchecked, would lead to unphysical results. While an extensive toolkit devoted to identifying, mitigating, and ultimately correcting such errors has been developed for Abelian groups, non-commuting symmetry operators complicate the implementation of similar schemes in non-Abelian theories. Here, we discuss two techniques for error mitigation through symmetry verification, tailored for non-Abelian lattice gauge theories implemented in noisy qudit hardware: dynamical post-selection (DPS), based on mid-circuit measurements without active feedback, and post-processed symmetry verification (PSV), which combines measurements of correlations between target observables and gauge transformations. We illustrate both approaches for the discrete non-Abelian group $D_{3}$ in 2+1 dimensions, explaining their usefulness for current NISQ devices even in the presence of fast fluctuating noise. Our results open new avenues for robust quantum simulation of non-Abelian gauge theories, for further development of error-mitigation techniques, and for measurement-based control methods in qudit platforms.
1 Introduction
@@
The rest of the paper is organized as follows: To make this work self-contained, in Sec. 2 we introduce the formalism needed to implement a quantum simulation of a non-Abelian LGT on qudit devices. In Sec. 3, we explain the general idea of the two techniques for ensuring non-Abelian gauge invariance. Section 4 specializes the concepts introduced so far to the $D_{3}$ model and presents our numerical analysis of the two protection schemes. Finally, we provide an overview of our work and discuss possible developments in Sec. 5. Further numerical results and details are presented in the Appendix.
2 Hamiltonian real-time evolution for discrete non-Abelian LGTs
@@
In this work, we are interested in a digital (Trotterized) time evolution of LGT models with a discrete non-Abelian symmetry group. While many of our concepts are general, we will focus on a pure gauge theory in 2+1 space-time dimensions in the Kogut–Susskind Hamiltonian formulation. We consider a square lattice with gauge variables residing on its links, see Fig. 1. To provide the necessary background, we review the basic formalism for quantum simulations of LGT models with a discrete non-Abelian group $G$, following the notation of Refs. [53, 54].
@@
Equation (2) can be verified by combining Eq. (2) with the representation of $U^{j}_{mn}$ in the computational basis {|$\{\ket{g}\}$.
@@
where $n_{t}=t/\Delta t$ is the number of Trotter steps. To efficiently encode groups of order $|G|>2$, we specialize our discussion to universal qudit devices  [36, 37, 38, 39, 40], where each fundamental unit of quantum information possesses more than two levels ($d>2$). This enables the implementation of physically local terms using single qudit gates  [58, 59]. The general framework developed for LGTs on qubit digital quantum simulators [60, 61, 54, 62] has already been expanded for qudit quantum hardware [59, 58, 63, 64, 65]. Furthermore, seminal proof-of-principle experiments on a truncated $U(1)$ group showcased the promising potential of a compact encoding of the gauge degrees of freedom [18]. In particular, qudit hardware is predicted to significantly reduce the circuit depth and increase the fidelity both for one-body [58] and multi-body terms [59] in the Hamiltonian. We summarize the essential ingredients and compare qubit and qudit-based implementations in Appendix A. For an efficient implementation, we assume the number of levels in the qudit to be at least as large as the group order, $d\geq|G|$. We discuss the resource advantage for this encoding of the gauge degrees of freedom, in contrast to a standard qubit approach, in Appendix C.
3 Verification of non-Abelian gauge symmetries
@@
3.1 Dynamical post-selection
@@
To further understand the result of this procedure, let us consider a situation where the incoherent error channel in Eq. (9) is described by probabilistically occurring unitaries. Such a scenario includes a broad range of errors, such as uncontrolled gate fluctuations, and will be at the basis of the noise model we use in our numerics in Sec. 4. Assuming the state at time $t$ satisfied the non-Abelian Gauss’ laws, we can approximate a noisy Trotter step of a single random (unitary) trajectory as
@@
3.2 Post-processed symmetry verification
@@
For a lattice gauge theory with symmetry group $G$, the symmetry-preserving subspace S is given by invariance under the gauge transformations $\Theta_{g,v}\ket{\psi}=\ket{\psi}$ for all group elements $g\in G$ and all of the $n_{\mathrm{v}}$ vertices $v\in\mathrm{V}$. In the case of a finite (Abelian or non-Abelian) $G$, the projector $\Pi_{\mathrm{s}}$ takes the form
@@
While this approach does not suffer from an exponentially decreasing number of valid shots, as does usual direct post-selection, the decompositions of Eq. (19) employ on the order of $|G|^{n_{\mathrm{v}}}$ operators. In Sec. 4.2.2, we discuss the resource requirements in more detail. Importantly, implementing this procedure does not require additional circuit elements in the form of ancillas or measurement gates and is not dependent on the commutation relations of the symmetry transformations.
4 Numerical benchmarks on the $D_{3}$ model
@@
We numerically simulate the system dynamics with exact diagonalization in the full Hilbert space of dimension $6^{L}$. The code is available at [70]. In our simulations, we use a fixed Trotter step $\Delta t/a=0.25$ and different noise strengths $\gamma$, and we set $\frac{1}{g^{2}}=0.5$ in order to work in a scenario where $H_{E}$ and $H_{B}$ compete.
4.1 Dynamical post-selection
@@
Figure 4: (a) Data clouds represent $\langle\mathcal{O}_{P1}\rangle$ at time $t/a$ on individual trajectories for DPS (red) and noisy (green) evolutions. The opacity reflects the fraction of trajectories retained after DPS. Solid lines represent ensemble averages. The DPS ensemble average follows very well the error-free evolution (black dots). In the inset, we show the standard deviation of DPS data (red dots) and the fraction $P_{\textrm{sur}}$ of trajectories that always preserved the gauge symmetries (black dots). (b-c) Gauge violations on vertex $v_{1}$ for the average of noisy evolutions without DPS (b) and for a single successful DPS trajectory (c). Reflection terms $g=3,4,5$ (solid lines) do not superimpose on a single noisy trajectory, while their averages do. Rotations are here represented by $g=2$ (dashed lines). Notably, the violation of a symmetry $\Theta_{g,v}$ can be partially suppressed also when other gauge charges are measured. The system is evolved with $\gamma=0.2$, $0.51/g^{2}=$ 0.5, and $\Delta t/a=0.25$ for $M=5000$ trajectories. The $x$-axis in panel (c) is zoomed in, compared to panels (a) and (b), to improve the readability.
@@
Here, $k_{g}$ is a normalization factor depending on $g$ such that $\textrm{GV}\in[0,1]$. The Gauss’ laws violation allows us to qualitatively compare post-selected and noisy dynamics and to assess the quality of the gauge-protection protocol. Figure 4(c) displays the dynamics of $\textrm{GV}_{g,v1}$ at the first vertex for a trajectory where the local symmetries are always successfully protected. Out of the six possible gauge transformations for each vertex, we can neglect the identity $g=0$ as it acts trivially. Moreover, we only monitor the violations corresponding to one rotation ($g=2$), since the $\mathbb{Z}_{3}$ subgroup of $\frac{2\pi}{3}$ rotations is cyclic, rendering conservation of the associated local symmetries equivalent. As shown in Fig. 4(c), the DPS protocol repeatedly suppresses the gauge violations through the auxiliary measurements, and their growth between two successive measurements of the same symmetry operator remains limited. This pattern remains constant even for longer times, although the noisy evolution becomes increasingly decoupled from the exact one. Moreover, gauge violations in Fig. 4(c) decrease even when other symmetries are monitored, which indicates that gauge violations of different local charges are correlated. The behaviour is markedly different in the bare noisy dynamics, shown in Fig. 4(b). Here, the ensemble-averaged violations increase until they reach a plateau, $0.5$ for reflections and $1/\sqrt{3}$ for rotations, corresponding to a fully mixed state [^2] The precise asymptotic value depends on the definition in Eq. (23), but in general every charge sector is uniformly populated after a characteristic time depending on $\gamma$..
@@
4.2 Comparing DPS and PSV
Let us now explicitly compare the error-mitigation capacity of the two protocols described in Sec. 3. First, we contrast the protection efficacy of both methods. After that, we discuss the relative scalings, which helps in assessing which method may be most suitable for a given platform.
4.2.1 Protection efficacy
@@ 4.2.2 Resource requirements
Additionally, one can estimate Eq. (19) through a random sampling of the group element combinations $\{g_{1},\dots,g_{n_{\mathrm{v}}}\}$. For every sample, one simultaneously measures the prescribed gauge transformation operator and its correlator with the gauge-invariant target observable (or cardinality-weighted cliques thereof). Analogous considerations for the qubit case [48, 47, 67] reveal such estimators to require a sampling overhead scaling as $\mathrm{Tr}[\Pi_{\mathrm{s}}\rho]^{-2}$. This overhead can generally be determined by analyzing the increase in the variance of the new estimator compared to the unmitigated, noisy one. The variance of a ratio of expectation values, such as those estimators, is then dominated quadratically by the inverse of the denominator. The decrease of $\mathrm{Tr}[\Pi_{\mathrm{s}}\rho]$ with circuit depth and system size is governed by the strength of (gauge-invariance breaking) error rates, and therefore the required sampling cost immediately benefits from any quantum hardware developments. In particular, for a given shot budget and error rate, one has a window of simulated time within which controlled sampling of the gauge-invariant expectation value is possible.
As a final remark, the feasibility and quality of the mid-circuit measurements are a fundamental requirement to implement DPS. While qudit hardware is still in a rapidly developing stage and the various platforms have yet to demonstrate their implementation, mid-circuit measurements are routinely used in qubit processors for a number of purposes, e.g., state preparation [22] or long-range entangling operations [71]. The extension to qudit devices requires some care, as the readout process becomes slower and more delicate because of the many levels involved, but faces no fundamental obstacle. In addition, efficient mid-circuit monitoring of ancilla qubits is a crucial building block for any error-correcting code and, as such, is actively pursued in many platforms aiming towards fault-tolerance.
5 Conclusions
@@
DPS is remarkably resilient against noise and the measurement strategy can be easily adapted to suit specific needs. For instance, the density of mid-circuit measurements can be tuned according to the system studied and the error channel expected on a specific hardware. The only requirement is that the measurement frequency must be above the quantum Zeno threshold for error mitigation in the system, otherwise, the system would enter an unprotected regime with random dynamics jumping between different gauge sectors [41]. The two-qudit gates used in the DPS circuit also contribute with an additional error rate. Nevertheless, as long as the number of additional two-qudit gates is sufficiently smaller than the ones already required for the Hamiltonian time evolution, this does not constitute a decisive limiting factor. On a qudit-based architecture, one needs $6(|G|-1)$ control-permutations per Trotter step per plaquette, using the strategy sketched in Appendix A. In contrast, checking one local symmetry $\Theta_{g,v}$ requires $4(|G|-n_{g})$ control-permutations, where $n_{g}$ is the number of states that transform trivially under the group element $g$. For the two-plaquette system considered above, this leads to 60 control-permutations for each Trotter step and 8 (10) for a reflection (rotation) symmetry check [^3] These number reflect the fact that each vertex has only three links in our geometry and the trivial subspace for the vertical links has dimension 4 for both reflection and rotations, see Appendix C.. This advantage shows a promising window of opportunity for future experimental studies on qudit-based hardware during the NISQ era. Moreover, the resource requirements can be mitigated by choosing an optimal subset of group elements for the symmetry measurements: gauge violations on a single vertex $v$ can be fully suppressed by monitoring only a suitably chosen generating set of $G$.
@@ Acknowledgements
Data availability statement
All numerical data and the code we used in this work are available at [70].
@@ Appendix A Trotterized time evolution of a discrete LGT
The unitary operator generated by the four-body interaction described by the magnetic term $H_{B}$ in Eq. (5) is more complicated to implement. A possible circuit decomposition [54], depicted in Fig. 6, is a generalization of a standard four-body Pauli-$Z$ interaction for qubits: three qudit control-multiplication gates acting on the fourth, onto which subsequently a trace gate is applied. Finally, the state of the four qudit registers is restored by applying the inverse of the multiplication gates. The procedure also requires the application of the inversion gate, which maps an element of the group $g$ into its inverse $\mathfrak{U}_{-1}\ket{g}=\ket{g^{-1}}$. The trace gate is diagonal in the group-element basis and implements the parametric operation
@@
The control-multiplication gate is defined as
@@
where the first qudit is the control and the second is the target. This is the most costly operation in quantum hardware, as it applies $|G|$ different operations (one of which, the identity, is trivial) depending on the state of the control qudit. The control-multiplication gate $\mathfrak{U}_{X}$ is the only two-qudit unitary required for the time evolution. To distinguish between $\mathfrak{U}_{X}$ and a group-multiplication controlled by specific group elements, we denote the latter as control-permutation. A single control-multiplication is thus a sequence of $|G|-1$ control-permutations.
On qubit hardware, in contrast, the QFT becomes a set of entangling gates. Its efficient implementation, in general, is not known, but a lot of effort has been made to find the optimal QFT for certain physically relevant non-Abelian groups [74]. One can gain some insight from $D_{4}$, whose QFT can be efficiently implemented using two entangling gates [75]. We can estimate that using qubits instead of qudits for $D_{4}$ would require $2\times L\times 2$ additional entangling gates for each Trotter step, where $L$ is the number of links in the lattice, the first factor $2$ is the number of entangling gates for every QFT, and the second factor $2$ represents the two QFTs we have to perform on every link at every Trotter step. In Ref. [59], the advantage of implementing the multiplication gate on qudits instead of qubits is analyzed for $\mathbb{Q}_{8}$, the quaternion group that is a non-Abelian subgroup of $SU(2)$ with cardinality $8$. In particular, they estimate a reachable fidelity of $99.6\%$ for qudits versus $21.4\%$ for qubits.